Circular No. 7932

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dw IN UGC 11376 D. Weisz and W. D. Li, University of California, report the LOTOSS

(cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT images taken on July 4.3 (mag ∼ 18.9) and 5.3 UT (mag ∼ 19.2). SN 2002dw is located at α = 18^{h}54^{m}15^{s}.25, δ = +48^{o}55′37′′.3 (equinox 2000.0), which is 8′′.1 east and 3′′.5 south of the nucleus of UGC 11376. Another (possibly cataclysmic) variable star (mag 19.7 on June 4.3; 19.4 on June 5.3), was found at position end figures 06^{s}.41, 18′′.5. A KAIT image taken on June 27.3 showed nothing at these two positions (limiting mag ∼ 20.0).

TRANSIENT X-RAY PULSARS R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; C. B. Markwardt, GSFC and University of Maryland; F. E. Marshall, GSFC; and S. Laycock and M. Coe, University of Southampton, report their detection of two transient x-ray pulsars in the direction of the Small Magellanic Cloud from RXTE Proportional Counter Array observations having 2^{o} FWZI fields-of-view centered at α = 0^{h}50^{m}00^{s}, δ = −73^{o}06′.0, and α = 0^{h}53^{m}53^{s}, δ = −72^{o}26′.7 (equinox 2000.0): “Pulsations at 82.4 ± 0.2 s from a previously unknown source were detected on June 19 and 26, and pulsations at 46.4 ± 0.1 s were detected on June 5 and 19. Source positions were determined from four scans, but accuracy is limited by the pulsations, which give variability on the scan timescales. For the 82.4-s pulsar, we find α(mean) = 0^{h}52^{m}55^{s} (standard deviation 2′) and δ(mean) = −72^{o}33′.0 (s.d. 8′), and so designate this source XTE J0052–725. The 46.4-s pulsar’s position is not near the scan centers and so is not well determined, but it is unlikely to coincide with the apparently inactive 46.6s pulsar 1WGA J0053.8–7226 (IAUC 6788, 6789). The intensity of XTE J0052–725 was ≈ 1.5 mCrab (2–10 keV) in all observations. The 46.4-s pulsar’s collimator-response-corrected intensity was poorly determined but was similar to that of XTE J0052–725 on June 19.”

COMET C/2002 C1 (IKEYA-ZHANG) Total visual magnitude estimates: May 21.08 UT, 5.4 (J. Bortle, Stor-

mville, NY, 10×50 binoculars); 31.96, 5.9 (M. V. Zanotta, Monte Tellero, Italy, 7×42 binoculars); June 8.54, 6.7 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 17.11, 7.1 (Bortle); 28.94, 8.4 (J. J. Gonzalez, Asturias, Spain, 11×80 binoculars); July 4.87, 8.9 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars).

2002 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green